For 550.0 mL of a buffer solution that is 0.150 M in CH3CH2NH2 and 0.140 M in CH3CH2NH3Cl, calculate the initial pH and the final pH after adding 0.010 mol of HCl.
Answer #1

this is a "basic" buffer, since there is weak base + conjugate acid:

CH3CH2NH2 + H2O <--> CH3CH2NH3+ + OH-

and

CH3CH2NH3Cl --> CH3CH2NH3+ + Cl-

so there is weak base and conjguate acid

now...

pOH = pKb + log(CH3CH2NH3+ / CH3CH2NH2 )

pKB for ethylamine = 3.19

so

initially:

mmol of base = MV = 550*0.15 = 82.5 mmol of base

mmol of conjugate = MV = 550*0.140 = 77 mmol of conjugate

calculate initial pH:

pOH = pKb + log(CH3CH2NH3+ / CH3CH2NH2 )

pOH = 3.19 + log(77 /82.5) = 3.16

pH = 14-3.16 = 10.84

after:

addition of 0.01 mol of HCl = 0.01*1000 = 10 mmol of H+

mmol of base = 82.5 -10 = 72.5 mmol

mmol of conjguate acid = 77 + 10 = 87

so

pOH = pKb + log(CH3CH2NH3+ / CH3CH2NH2 )

pOH = 3.19 + log(87/72.5) = 3.26918

pH = 14-3.26918= 10.730